CURRICULUM VITAE

1 PERSONAL DETAILS

Full name  Tomasz Aleksander Kowalewski
Date of Birth  25th September, 1947
Place of Birth  Łódź, Poland
Citizenship  Polish
Married with Teresa Ryczerz  1969
Children  Ingrid 1970 - 2004, son Oskar 1973

Work Address Institute of Fundamental Technological Research Polish Academy of Science (IPPT PAN) Świȩtokrzyska 21, PL 00-049 Warszawa, Poland Phone: +48-228269803, Fax: +48-228274699 e-mail: tkowale@ippt.gov.pl http://www.ippt.gov.pl/˜ tkowale

Home Address Sikorskiego 11 m 81, PL 02-758 Warszawa Phone: +48-224074330, +48-602193472 (mobile)

2 ACADEMIC QUALIFICATIONS

Degrees M.Sc. in Physics with Dean’s Distinction Award, Warsaw University, 1969. Ph.D., Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, 1982. Habilitation, Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, 1995.

3 APPOINTMENTS

Present Professor, head of the Department of Mechanics and Physics of Fluids, Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Warsaw, Poland.

June 1993 – June 1994 Wissenschaftlicher Mitarbeiter (Senior Research Associate), Max-Planck-Institut für Strömungsforschung, Göttingen, Germany March 1988 – May 1993 Wissenschaftlicher Mitarbeiter (Senior Research Associate), Institut für Angewandte Mechanik und Strömungsphysik, Georg-August-Universität, Göttingen.

Oct. 1985 – Feb. 1988 Visiting Scholar, research fellowship by Max-Planck-Institut, Göttingen.

Sept. 1969 – June 1987 Asystent (Research Assistant) and Adiunkt (Senior Research Associate), Institute of Fundamental Technological Research, Polish Academy of Science (IPPT PAN), Department of Fluid Mechanics, Warsaw, Poland.

4 SCHOLARSHIPS

Aug. – Sept. 1978 Visiting scholar at the Delft University, Holland (by Prof. J.A. Steketee). Hot-Wire Anemometry.

Nov. – Dec. 1980 Visiting scholar at the Max-Planck-Institut für Strömungsforschung, Göttingen (by Prof. E.-A. Müller). Investigations of twophase flow with help of Laser-Doppler-Anemometry.

Mar. – June 1982 Visiting scholar at the Stanford University (by Prof. A. Acrivos). M. Curie-Sklodowska joined Polish-American project on sedimentation and flow of suspensions.

Oct. 1985 – Feb. 1988 Scientific fellowship from the Max-Planck-Society, Institut für Strömungsforschung, Göttingen (by Prof. E.-A. Müller). Experimental modeling of fuel injection and evaporation; fuel film deposition on the walls of engine manifolds.

Feb. – Apr. 1993 Visiting Fellow at the University of New South Wales, School of Mechanical and Manufacturing Engineering. (by Prof. G. de Vahl Davis). Experimental and numerical investigation of the natural convection.

Nov. 1993 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. A. Yarin), Haifa, Israel. Investigations of liquid jet instability.

Mar. 1995 – Jan. 1996 Visiting Fellow at the Israel Institute of Technology, Lady Davis Fellowship, TECHNION, Dept. of Aerospace Engineering (by Prof. M. Wolfshtein), Haifa, Israel. Investigations of droplet combustion. Teaching (two semesters course) Advanced Methods in Aerodynamics.

April 1997 Visiting Professor at the University Paris XI (Orsay), lecturing: “Advanced experimental methods in fluid mechanics”.

April-May 1998 Visiting Professor at the University “La Sapienza” and Roma Tre (Rome), application of the PIV method for combustion flow

Nov. 1998 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

March 1999  Visiting Professor at the University Paris XI (Orsay)

Dec 1999 - Jan 2000, Sept. - Nov. 2000, Dec. 2001 - Jan 2002

Visiting Scholar at Arizona State University

June 2000 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

October 2001 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Mechanical Engineering (by Prof. G. Hetsroni), Haifa, Israel. Joint project on turbulent flow in the channel

March 2002  Visiting Professor at the University Paris XI (Orsay)

Nov. 2002 Short visit at the Israel Institute of Technology, TECHNION, Dept. of Aerospace Engineering (by Prof. Y. Levy), Haifa, Israel. Joint 5FP EU project on combustion FLOXCOM

March 2003 Visiting Professor at the University of New South Wales (Sydney)

5 RESEARCH EXPERIENCE

1969 – 1985 Research work at the Institute of Fundamental Technological Research (Warsaw), during the experimental study towards Ph.D. degree on low Reynolds number flow of suspensions; experimental investigations of the jet break-up, modeling of the fibre spinning.

1986 – 1987 Research work at the Max-Planck-Institut, Evaporation of fuels and retrograde fluids; atomisation of fuels; investigation of natural convection.

1988 – 1991 Research grants from Deutsche Forschungsgemeinschaft through the University Göttingen: Experimental investigations of the instabilities of the evaporating liquid jets (DFG 201/18) and

1991 – 1993  Dynamic properties of the jets of liquid mixtures (DFG201/20).

1994 – 1997 Research grant from Polish Scientific Committee (KBN): Flow pattern selection for natural convection with phase change in a lid cooled cavity (KBN 2P404000107)

1995 – 1996 9 months Lady Davis Fellowship, Application of PIV and image processing for diagnostics of flow with combustion, Israel Institute of Technology, TECHNION, Haifa

1996 – Research collaboration (project conjoins) IPPT PAN – LIMSI CNRS, L’etude numerique et experimentele de la dynamique du gossissement et du decrochement des bulles de vapeur dans le cas de l’ebullition heterogene

1998 – 2002 Research collaboration IPPT PAN - TECHNION in framework of Polish Academy of Sciences – Israeli Academy of Sciences agreement, on Turbulent thermal heat transfer in a channel flow Other activities Nonlinear dynamic of the oscillating droplet; evaporation of droplets; evaporation induced surface instabilities; image processing; high-Speed video imaging; flow visualisation; tracking of tracer particles in the flow; natural convection; convection with phase changes; Particle Image Velocimetry (PIV), liquid crystals thermography (PIT), computational fluid dynamics, micro-fluidics.

6 MAIN RESEARCH TOPICS

Keywords: Fluid mechanics, micro-fluidics, multiphase flow, free surface flow, thermally driven flow, flow with phase changes. Experimental and numerical methods.

• Experimental investigations of liquid jet break-up. Evidence of “super” thin liquid bridges between liquid jet and the droplet. Instability of evaporating jets, modeling of droplet and liquid jet evaporation. Evaporation of emulsions [9, 10, 13, 31, 35, 43-45].

• Experimental and numerical investigation of nonlinear droplet oscillations. Description of droplet deformation. Development of non-intrusive method to measure dynamic surface tension, measurements of the physical parameters of evaporating surface by oscillating droplet method [11, 18, 20, 21, 25, 26, 36, 37, 40, 42].

• Development of high speed video recording methods based on CCD “frame-transfer” camera and image processor. Application of the method for investigation of droplets and jet instabilities [8, 11, 15, 19, 32].

• Experimental and numerical investigation of natural convection, identification of new flow structures; investigation of natural convection in presence of phase change (freezing) [12, 14, 16, 24, 27, 29, 41, 46-54].

• Development of new experimental technique for instantaneous measurement of temperature and velocity two-diemnsional fields in thermal flow, based on application of liquid crystal as tracers [7, 14, 29, 34,48,49].

• Investigations of low Reynolds number two-phase flow in tubes (solid and liquid suspensions). Application of new measuring techniques for measuring velocity and particle concentration, laser doppler and ultrasound doppler anemometry [1-6].

7 EDUCATIONAL EXPERIENCE

Direct supervision of 7 M.Sc. and 3 Ph.D. students graduated at the Georg-August University Göttingen; 2 Ph.D. and 4 M.Sc. students at Warsaw University of Technology, 4 Ph.D students at IPPT PAN.

Graduate course: Advanced Methods in Aerodynamics (two semesters) given at TECHNION, Dept. of Aerospace Engineering.

Languages: English, German - good; Russian, French - weak.

8 REVIEWER

Journals

1. Journal of Fluid Mechanics

2. Physics of Fluids

3. Archives of Mechanics

4. Computer Assisted Mechanics and Engineering Sciences

5. Journal of Engineering Mathematics (member of Board of Assoc. Referees)

6. Journal of Heat Mass Transfer

7. Applied Thermal Engineering

8. JMEMS

9. Meccanica

10. APS PRL

11. APS PRE

12. AIChEJ 13 Machine Graphics Vision

14. Fluid Dynamics & Materials Processing (member of Editorial Board)

15. Bulleting of Polish Academy of Sciences (member of Editorial Board)

16. Int. J. Thermal Scs.

17. Industrial & Engineering Chemistry Research

18. Exp. in Fluids

19. J. Micromechanics and Microengineering

Research grants, professor nomination

1. National Science Foundation, Grants Committee, Washington, USA

2. Australian Research Council, Grants Committee, Canberra, Australia

3. The University of New South Wales, Promotion to Professor, Sydney, Australia

4. Deutsche Forschungsgemeinschaft (DFG), Grants Committee, Germany.

5. Israel Science Foundation (ISF), Grants Committee, Israel.

6. State Scientific Committee (KBN), Grants Committee, Poland.

9 HONOURS

Scientific Secretary, XVIIth Symposium on Advanced Problems and Methods in Fluid Mechanics, Sobieszewo (Poland), Sept.1985.

Scientific Secretary, Publication Subcommittee, IUTAM Symposium on Adiabatic Waves in Liquid-Vapour Systems, Göttingen (Germany), Aug. 1989.

Invited paper at 2nd European Fluid Mechanics Conference, Warsaw, September 1994

Invited paper at Euromech Colloquium 355 on Interfacial Instabilities, Paris, September 1996,

The Polish Academy of Sciences Award (Nagroda Wydzialu IV PAN) for research on jets and droplets, December 1996. International Symposium on Advances in Computational Heat Transfer CHT97, Cesme 1997, memeber of the Scientific Committee and the expository review lecture. European Science Foundation project on Applied Mathematics for Industrial Flow Problems, Polish memeber of the Steering Committee, memeber of the Scientific Committee, “International AMIF conferences”, 1998, 2000, 2002, International Conference on Computational Heat and Mass Transfer, Cyprus, April 1999, memeber of the Scientific Committee. Invited paper at FLUVISU 99, 8me colloque nationale de visualisation et de traitment d’images en mecanique des fluides, Toulose, June 1999. Chairman of the Euromech Colloquium 406 “Image processing methods in applied mechanics”, Warsaw, May 1999. Chairman of the European Science Foundation AMIF Workshop PCC99 “Phase Change with Convection”, Warsaw, June 1999. Memeber of the International Board “9th and 10th Int. Flow Visualization Symposium”, 2000, 2002, Memeber of the Scientific Committee, “International Symposium on Advances in Computational Heat Transfer CHT01”, Australia 2001. Memeber of the Scientific Committee and invited paper at “International Symposium on Imaging and Visualization in Transport Phenomena VIM-01”, Antalya, 2002. Silver Cross Award by President of Poland for scienitifc achievements, 2002 Coordinator of “Phase Change with Convection”, CISM Advanced School, Udine, 2002. Secretary General, “International Congress of Theoretical and Applied Mechanics”, Warsaw, 2004. Member of the XCCC IUTAM, 2004-2008.

10 PRESENTATIONS AT INTERNATIONAL CONFERENCES

September 1977 XIII Biennale on Adv. Fluid Mechanics  Kortowo, Poland
September 1980 Flow Visualization II  Bochum, Germany
September 1981 XV Biennale on Adv. Fluid Mechanics  Jahranka, Poland
August 1984  16. IUTAM Congress  Lyngby, Denmark
April 1985  Euromech Colloquium  Cambridge, UK
April 1986  Taylor’s 100 Birth. Aniv. Symp.  Cambridge, UK
August 1986  Flow Visualization IV  Paris, France
November 1987 ICALEO  San Diego, USA
April 1988  GAMM  Wien, Austria
August 1988  17. IUTAM Congress  Grenoble, France
August 1989  IUTAM Symposium on Topol. Fluid. Mech.  Cambridge, UK
September 1989 XIX Biennial on Adv. Fluid Mechanics  Kozubnik, Poland
December 1989 10 Aust. Fluid Mech. Conf.  Melbourne, Australia
August 1991 1. Euromech Fluid Mech. Conf. Cambridge, UK
December 1991 11. ABCM Mech. Eng. Conf.  San Paulo, Brazil
August 1992  18. IUTAM Congress  Haifa, Izrael
October 1992  Flow Visualization VI  Yokohama, Japan
October 1993  1. Japan-Polish Joint Seminar  Tokyo, Japan
June 1994  2. Japan-Polish Joint Seminar  Pułtusk, Poland
August 1994  10th Int. Heat Transfer Conference  Brighton, UK
September 1994 2. Euromech Fluid Mech. Conf.  Warszawa, Poland
September 1994 Jap.-Centr. Europe Workshop  Pułtusk, Poland
June 1995  Euromech Colloquium 335  Rome, Italy
July 1995  3 ICIAM Congress  Hamburg, Germany
August 1996  19. IUTAM Congress  Kyoto, Japan
September 1996 Euromech Colloquium 355  Paris, France
December 1996 Heat’96  Kielce, Poland
May 1997 Advances in Computational Heat Transfer Cesme, Turkey
September 1997 3. Euromech Fluid Mech. Conf.  Göttingen , Germany
June 1998  Advanced Computational Methods in Heat Transfer Cracow , Poland
September 1998 8th Flow Visualization  Sorrento, Italy
October 1998  AMIF - ESF Conference  San Feliu, Spain
May 1999  Euromech 406 Colloquium  Warszawa, Poland
June 1999  FLUVISU 99  Toulouse, France
June 1999  PCC99 AMIF-ESF Workshop  Warszawa, Poland
August 1999  Cold-Region Thermal Eng. ISTESCR’99  Darmstadt, Germany
September 1999 Laser Anemometry Adv. EALA’99  Rome, Italy
August 2000  9th Flow Visualization  Edinburg, Scotland
August 2000  20. IUTAM Congress  Chicago, USA
June 2001  CMEM 2001  Alicante, Spain
December 2001 3d ISEH  Tempe (AZ), USA
May 2002  VIM’01  Belek, Turkey
June 2002  HEAT2002  Baranow, Poland
August 2002  12 Heat Transfer Conf.  Grenoble, France
October 2002  Eurotherm 71  Reims, France
June 2003  Eurotherm 69  Bistra, Slovenia
August 2003  5. Euromech Fluid Mech. Conf.  Tuluse, France
August 2004  ICTAM04  Warsaw, Poland

May 2005  4th ICCHMT  Paris-Cachan, France
September 2005 CDMM2005  Warsaw, Poland
June 2006 4th ASME ICNMM Limerick, Irleand
June 2006  6. Euromech Fluid Mech. Conf. Stockholm, Sweden
September 2006 12th Flow Visualization  Goettingen, Germany

11 PUBLICATIONS

1. Experimental determination of filtration coefficients along bunch of fibres (in Polish), with W. Kalita, IFTR Reports - Prace IPPT, vol. 28/1976, Warszawa 1976.

2. Optical metod for measuring concentration profiles of droplet suspension flowing through a tube (in Polish), IFTR Reports - Prace IPPT, vol. 49/1976, Warszawa 1976.

3. Velocity profiles of suspension flowing through a tube, Arch. Mech., 32, pp. 857-865, 1980.

4. Laser Doppler measurements of the velocity profiles of concentrated droplet suspensions flowing through a tube, Bericht 118/1980, Max-Planck-Institut für Strömungsforschung, Göttingen 1980.

5. Experimental investigation of laminar flow of suspension in channels (in Polish), Ph.D. dissertation, IPPT PAN, Warsaw 1982.

6. Laser method of particles distribution measurements in the flow of suspensions, in: Flow Visualization II, Bochum 1980, Ed. W. Merzkirch, Hemisphere 1982, pp. 541-546.

7. Concentration and velocity measurements in the flow of droplet suspensions, Exp. in Fluids, 2, pp. 213-219, 1984.

8. An experimental study of the lateral migration of a droplet in a creeping flow, with W. Hiller, Exp.in Fluids, 5, pp. 43-48, 1987.

9. Simultaneous measurement of temperature and velocity fields in thermal convective flows, with W. Hiller, in: Flow Visualization IV, Paris 1986, Ed. Claude Veret, Hemisphere 1987, pp. 617-622.

10. Eine einfache Hochgeschwindigkeitskamera mit CCD- Sensor, with W. Hiller, Bericht 8/1987, Max-Planck- Institut für Strömungsforschung, Göttingen 1987.

11. Freistrahlversuche zum genauen Einsatz der Verdampfungseffekte bei Einspritzung von verschiedenen motorischen Kraftstoffen, with H. Chaves, H.-D. Speckmann and G.E.A. Meier, Bericht 102/1987, Max-Planck-Institut für Strömungsforschung, Göttingen 1987.

12. Untersuchung des Verdampfungsverhaltens von Einspritzstrahlen in einem Saugrohrmodell, with H. Chaves and G.E.A. Meier, Bericht 103/1987, Max-Planck-Institut für Strömungsforschung, Göttingen 1987.

13. An optical method for surface tension measurements of dispersed liquid droplets, with W.J. Hiller, in: Proc. of ICALEO’87, 6th Int. Congress on Application of Laser and Electro-Optics, Optical Methods in Flow and Particle Diagnostics, vol. 63. Ed. W. Stevenson, Publ. Laser Institut of America, pp. 106-110, Toledo 1988.

14. Three-dimensional structures in laminar natural convection in a cubic enclosure, with W.J. Hiller, St. Koch, Proc. of 1st. Word Conference on Exp. Heat Transfer, Fluid Mech., and Thermodyn., Dubrovnik 1988, ed. R.K. Shah et al., Elsevier Scis. Publ., New York, Amsterdam, London, pp. 722-729, 1988.

15. Similarity in the behaviour of initially saturated or subcooled liquid jets discharging through a nozzle, with H. Chaves, Th. Kurschat, G.E.A. Meier, E.A. Müller, Chem. Phys. 126, pp. 137-143, 1988.

16. Simultane Erfassung von Temperatur- und Geschwindigkeitsfeldern in einer thermischen Konvektionsströmung mit ungekapselten Flüssigkristalltracern, with W.J. Hiller, St. Koch, in: 2D-Messtechnik DGLR-Workshop 18-19. Oct. 1988, Markdorf, DGLR-Bericht 88-04, pp. 31 - 39, Bonn 1988.

17. Application of the frame transfer charge-coupled device for high speed imaging, with W.J. Hiller, Optical Eng. 28 pp. 197-200, 1989; also in: Proc. 18th Int. Congr. of High Speed Photography and Photonics, 28. Aug.-2. Sept. 1988 Xian China, SPIE Vol. 1032, pp. 763 - 765, Washington 1990.

18. Three-dimensional structures in laminar natural convection in a cubic enclosure, with W.J. Hiller, St. Koch, Exp. Thermal and Fluid Sci., 2, pp. 34-44, 1989.

19. Surface tension measurements by the oscillating droplet method, with W.J. Hiller, Physico-Chem. Hydrodynam., 11, pp. 103-112, 1989.

20. Optical investigation of oscillating liquid droplets, with W.J. Hiller, ZAMM 69(6), pp. 629-630, 1989.

21. Schnelle Bildaufzeichnung mit CCD-Kameras und gepulsten LEDs - High Speed Image Recording, with W.J. Hiller, B. Stasicki, Laser und Optoelektronic 21, pp. 64-67, 1989.

22. Experimental analysis of free oscillating liquid drops, with W.J. Hiller, in: Proceedings of the 10th Australasian Fluid Mechanics Conference, University of Melbourne, Dec. 1989, pp. 7.21 -7.24, 1989.

23. Liquid microjets - a useful tool for the measurement of material properties, with W.J. Hiller, in: Proc. of the 10th ABCM Mechanical Engineering Conf., Rio de Janeiro (Brasil), Dec. 1989, Eds. M. Hirata et al., pp. 423 - 426, COPPE/UFRJ 1989.

24. Vergleichende Grunlagenstudie über das Schwingungsverhalten osczillierender Wurzelkanalinstrumente, with Griesinger H.R., W.J. Hiller, Kreter F., Zahnärztliche Praxis 5, pp. 168-173, 1989.

25. Vibrationsinduzierte Strömungsfelder und Kavitationseffekte dreier Systeme zur Wurzelkanalaufbereitung im Modellversuch, with Griesinger H.R., W.J. Hiller, Kreter F., Zahnärztliche Praxis 6, pp. 213-217, 1989.

26. Experimental and Numerical investigation of natural convection in a cube with two heated side walls, with W.J. Hiller, St. Koch, G. de Vahl Davis, M. Behnia, in: Proc. of IUTAM symposium on topological fluid mechanics, Cambridge U.K., Aug. 13-18, 1989, pp. 717 - 726, CUP 1990.

27. Experimental and theoretical investigations of large amplitude oscillations of liquid droplets, with E. Becker, W.J. Hiller, J. Fluid Mech. 231, pp. 189-210, 1991 28. Unsteady droplet evaporation with W.J. Hiller, in: Proc. of 11th ABCM Mech. Eng. Conf. São

Paulo (Brasil), Dec. 1991, Ed. Esp. da Revista Brasileira de Ciências Mecânicas São Paulo, pp. 17-20, 1991.

29. Visualization of 3-D natural convection - comparison with numerical results, with W.J. Hiller, St. Koch, K. Range, M. Behnia, G. de Vahl Davis, in: Proc. of 11th ABCM Mech. Eng. Conf. São Paulo (Brasil), Dec. 1991, Ed. Esp. da Revista Brasileira de Ciências Mecânicas São Paulo, pp. 21-24, 1991.

30. Computer-aided discrimination of slow and fast tracer paths, with H.-H. Bartels-Lehnhoff, W.J. Hiller, Exp. in Fluids, 13, pp. 239-248, 1992; also in: Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992, Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 873–877, 1992.

31. Visualisation of 3-D natural convection, with W.J. Hiller, St. Koch, P. Mitgau, K. Range, Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992 Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 674–678, 1992.

32. Charge-coupled devices in flow visualisation, with W.J. Hiller, V. Llorach Forner, B. Stückrad, M. Behnia, in: Proceedings of The Sixth International Symposium on Flow Visualisation, Yokohama 1992, Eds. Tanida Y. & Miyashiro H., Springer-Verlag, pp. 695–699, 1992.

33. Behaviour of small diameter evaporating jets, with W.J. Hiller, Behnia M.,, in: Proceedings of the 11th Australasian Fluid Mechanics Conference, University of Tasmania, Hobart, Australia, 14-18. Dec. 1992, pp. 905–908, U.T. 1992.

34. High speed frame transfer CCD, with W.J. Hiller, Tatarczyk Th., in: Proc. 20th Int. Congr. of High Speed Photography and Photonics, 21-25. Sept. 1992 Victoria, Canada, SPIE Vol. 1801, pp. 595–601, Washington 1993.

35. Wake patterns of a piston gliding in a rotating circular duct, with W.J. Hiller, in: Bluff-Body Wakes, Dynamics and Instabilities, IUTAM Symposium, Göttingen, Germany, Sept. 7-11, 1992, Eds. H. Eckelmann et al., Springer-Verlag, pp. 165–169, 1993.

36. Onset of natural convection in a cube, with W.J. Hiller, St. Koch, F. Stella, Int. J. Heat Mass Transfer, 13, pp. 3251–3263, 1993.

37. An experimental study of evaporating small diameter jets, with W.J. Hiller, M. Behnia, Physics of Fluids A 5, pp. 1883–1890, 1993.

38. Measurement of dynamic surface tension by the oscillating droplet method, with B. Stückrad, W.J. Hiller, Exp. in Fluids, 15, pp. 332-340, 1993.

39. Nonlinear dynamics of viscous droplets, with E. Becker, W.J. Hiller, J. Fluid Mech., 258, pp. 191-216, 1994.

40. Computational and experimental visualisation in heat and mass transfer problems, with W.J. Hiller and G. de Vahl Davis, in Proc. of the First Japanese-Polish Joint Seminar in Advanced Computer Simulation, Tokyo, Nov. 8-9, 1993, pp. 60–69, Edts. Akiyama, Kleiber, Wolański, University of Tokyo Feb. 1994.

41. Verfolgung von Teilchen in einer dreidimensionalen Strömung, with P.M. Mitgau, W.J. Hiller, ZAMM 74(5), pp. T394–396, 1994.

42. Reduction of nonlinear dynamic systems by phase space analysis, with E. Becker, U. Brosa, Computer Assisted Mechanics and Eng. Sci., 1, pp. 39-48, 1994.

43. Experimental and numerical study of three-dimensional natural convection and freezing in water, with C. Abegg, G. de Vahl Davis, W.J. Hiller, St. Koch, E. Leonardi, G.H. Yeoh, Proc. of 10th International Heat Transfer Conference, Brighton, England, Edt. G.F. Hewitt, vol.4, pp. 1-6, IChemE 1994.

44. Nonlinear Oscillations of Viscous Droplets, with D. Bruhn, Proc. of Japanese-Centr. European Workshop on Adv. Comp. in Eng., Pułtusk 1994, pp. 63-68, Edts. Akiyama, Kleiber, IPPT PAN Warszawa 1994.

45. Experimental research on briquette destruction caused by refraction waves, with M. Gawor, J. Rysz, A. Smolarski, Archives of Mining Scs., vol. 39, pp. 313-330, 1994.

46. Selected free surface flow problems - liquid jets and drops, (in Polish) Habilitation, Prace IPPT PAN 3/1995, pp.1-98, Warszawa 1995.

47. Experimental investigations of oscillating liquid droplet, with D. Bruhn, W.J. Hiller, F. Obermeier, ZAMM vol. 76/5, pp.77-78, 1996.

48. On the separation of droplets from a liquid jet, Fluid Dynamics Research 17, pp. 121-145, 1996.

49. Distribution of particles suspended in 3-d laminar convection flow, with A. Yarin, W.J. Hiller, St. Koch, Physics of Fluids 8, pp. 1130-1140, 1996.

50. Experimental and numerical investigations of natural convection in freezing water, with A. Cybulski, in Int. Conf. on Heat Transfer with Change of Phase, Mechanics vol. 61 p.II, pp. 7-16, Kielce 1996.

51. Experimental validation of numerical codes in thermally driven flows, Prace IPPT PAN 4/1997, pp.1-20, Warszawa 1997.

52. Natural convection with phase change (in Polish), with A. Cybulski, Prace IPPT PAN 8/1997, pp.1-58, Warszawa 1997.

53. Estimation of acoustic streaming: theoretical model, Doppler measurements and optical visualisation, with A. Nowicki, W. Secomski, J. Wojcik, European J. of Ultrasound 7, pp. 73-81, 1998.

54. Experimental validation of numerical codes in thermally driven flows, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.1-15, New York 1998.

55. An experimental benchmark for freezing water in the cubic cavity with M. Rebow, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.149-156, New York 1998.

56. Fixed grid finite elemnt analysis of solidification with J. Banaszek, M. Rebow, in Adv. in Computational Heat Transfer, G. de Vahl Davis, E. Leonardi (eds), Begel House Inc., pp.471-478, New York 1998.

57. Free surface natural convection in a differentially heated rectangular cavity, with G. de Vahl Davis, E. Leonardi, in Advanced Computational Methods in Heat Transfer V, A.J. Nowak, C.A. Brebbia, R. Bialecki, M. Zerroukat (eds.), pp. 145-154, CMP 1998.

58. Particle Image Velocimetry with Optical Flow, with G. M. Quénot, J. Pakleza, Experiments in Fluids 25, pp.177-189, 1998.

59. Analysis of validation of the semi-implicit FEM algorithm for natural convection and freezing water, with J. Banszek, M. Rebow, K. Blogowska, in Computational Technologies for Fluid/Thermal/Structural/Chemical Systems with Industrial Applications, V.V. Kudriavtsev, W. Cheng (eds), ASME PVP-Vol.377-1, pp.119-126, 1998.

60. Fluid Image Velocimetry for Unseeded flow, with Y. Levy, B. Golovanevsky in Proc. of 9th Int. Symp. on Application of Laser Technique in Fluid Mech., July 13-16, Lisbon, vol. II, pp. 34.7.1-34.7.9, 1998.

61. Characterisation of the interaction between a boundary layer and a cavity using Digital Particle Velocimetry with Optical Flow, with G. Quenot, A. Rambert, P. Gougat, in Proc. of 9th Int. Symp. on Application of Laser Technique in Fluid Mech., July 13-16, Lisbon, vol. II, pp.39.4.1-39.4.7, 1998.

62. Study of free convection in binary systems with phase change, with J. Banaszek, M. Rebow, P. Furmanski, T. Wisniewski, in Heat Transfer 1998, Kyongju, Korea, edt. J.S. Lee, vol.7, pp. 211-216, Taylor and Francis Inc., 1998.

63. Particle Image Velocimetry and Thermometry in Freezing Water, with A. Cybulski, M. Rebow, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109 , pp. 24.1-24.8, Edinburgh, 1998.

64. Particle Image Velocimetry using Optical Flow for Image Analysis, with G. Quenote, J. Pakleza, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109, pp. 47.1-47.11, Edinburgh, 1998.

65. Infrared and Liquid Crystal Thermography in Natural Convection, with T. Wisniewski, M. Rebow, K. Blogowska, in 8th Int. Symposium on Flow Visualization, Sorrento, eds. G.M. Carlomagno and I. Grant, CD ROM Proceedings ISBN 0953399109, pp. 212.1-212.8, Edinburgh, 1998.

66. Thermochromic liquid crystals in heat transfer research, with J. Stasiek, invited paper at Heat Mass Transfer Conference, Swieradow, Sept. 1998, Prace Polit. Wroclaw. 53/9, vol. 1, pp. 401-424, Wroclaw 1998.

67. Freezing of water in the differentially heated cubic cavity, with M. Rebow, Int. J. of Comp. Fluid Dyn. vol. 11, pp. 193-210, 1999.

68. Particle Image Velocimetry and Thermometry in Freezing Water, with A. Cybulski, frontispiece (colour page) in J. of Visualization, vol. 1, p. 335 , 1999.

69. Effect of finite wall conductivity on flow structures in natural convection, with E. Leonardi, V. Timchenko, G. de Vahl Davis, paper at CHMT99, Proceedings of the International Conference on Computational Heat and Mass Transfer, Cyprus 26-29 April, 1999, eds. A.A. Mohamad & I. Sezai, pp. 182-188, Eastern Mediterranean University Printinghouse 1999.

70. Particle Image Velocimetry and Thermometry Using Liquid Crystals, invited paper at FLUVISU 99, 8me colloque nationale de visualisation et de traitment d’images en mecanique des fluides, Toulose, June 1-4, 1999, ENSICA, pp. 33-48, Image Service 1999.

71. Measurement of Temperature and Velocity of Freezing Water using Liquid Crystal Tracers in Advances in Cold-Region Thermal Engineering and Sciences, eds. K. Hutter, Y. Wang and H. Beer, Lecture Notes in Physics vol. 533, pp. 83-90, Springer Verlag, Berlin 1999.

72. Particle Image Velocimetry for Vapour Bubble Growth Analysis with J. Pakleza, A. Cybulski; Proceedings of the 8th International Conference on Laser Anemometry, pp. 243-250, Univ. of Rome La Sapienza, 1999.

73. An axisymmetry-breaking instability of axially symmetric natural convection with A.Yu. Gelfgat, P.Z. Bar-Yoseph, A. Solan; Int. J. Trans. Phenomena: vol.1, pp. 173-190, 1999.

74. Semi-implicit FEM analysis of natural convection in freezing water with J. Banaszek, Y. Jaluria, M. Rebow; Num. Heat Transfer, Part A, vol. 36, pp. 449-472, 1999.

75. Phase-change problems with free convection: fixed grid simulation with M. Giangi, F. Stella; Comp. & Vis. in Scs. vol.2 (2/3), pp. 123-130, 1999.

76. Large-scale structures forming in a cross flow: particle image velocimetry conditional analysis with R. Camussi, A. Stella, G. Guj; Machine Graphics & Vision vol. 8 (4), pp. 579-595, 1999.

77. Natural convection during ice formation: numerical simulation vs. experimental results with M. Giangi, F. Stella, E. Leonardi; CAMES, vol 7(3), pp. 321-342, 2000.

78. Natural convection with phase change in one-component and binary systems (in Polish), with J. Banaszek, P. Furmanski, M. rebow, A. Cybulski, T.S. Wisniewski; Prace IPPT PAN 3/2000, pp.1-132, Warszawa 2000.

79. Tracking of thermal structures from infrared camera by PIV method, with G. Hetsroni, B. Hu, A. Mosyak in 9th Int. Symposium on Flow Visualization, Edinburgh, Edt. I. Grant, CD ROM Proceedings, pp. 131.1-9, Edinburgh, 2000.

80. Visualization of vapour bubble growth, with J. Pakleza, J.-B. Chalfen, M.-C. Duluc, A. Cybulski in 9th Int. Symposium on Flow Visualization, Edinburgh, Edt. I. Grant, CD ROM Proceedings, pp. 176.1-9, Edinburgh, 2000.

81. Tracking of coherent thermal structures on a heated wall by means of IR thermography with G. Hetsroni, B. Hu, A. Mosyak; Experiments in Fluids, vol. 30, pp. 286-294, 2001.

82. Experimental model for casting problems, with A. Cybulski, T. Sobiecki, in Computational Methods and Experimental Measurements, Eds. Y. V. Esteve, G.M. Carlomagno, C.A. Brebia, pp. 179-188, WIT Press, Southampton 2001.

83. Experimental investigations on thermal, thermocapillary, and forced convection in Czochralski crystal growth configuration, with J. Aleksic, J.A. Szymczyk, A. Leder in Computational Methods and Experimental Measurements, Eds. Y. V. Esteve, G.M. Carlomagno, C.A. Brebia, pp. 627-636, WIT Press, Southampton 2001.

84. Visualization of Vapour Bubble Growth (fontipiece), with J. Pakleza, J.-B. Chalfen, M.-C. Duluc, A. Cybulski, J. of Visualization, vol. 4, p. 212, 2001

85. Methodologie experimentale pour l’etude de la croissance de bulles de vapour en site isole, with J. Pakleza, J.-B. Chalfen, M.-C. Duluc, A. Cybulski, in Science et technologies thermiques face au XXIe siecle, SFT2001, Acte du congres annuel Nantes, Edts. C. Castelain, D. Delaunay, pp. 467-472, Elsevier 2001,

86. Particle Image Velocimetry and Thermometry using Liquid Crystals Tracers, PIV’01 4th Int. Symp. on Particle Image Velocimetry CD ROM Proc, pp.1134.1-10, Edt. J. Kompemhans, DLR Mitteilung 2001-03, Goettingen 2001.

87. Vertical transport ond mixing in the Salt-Lake basin, with P. Monti, M. Princevac, W.C. Chan, H.J. Fernando, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00286.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

88. Slope flow measurements during vertical transport and mixing (VTMX) field experiment, Salt Lake City 2000, with M. Princevac, H.J. Fernando, P. Monti, W.C. Chan, J. Anderson, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00238.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

89. Application of liquid crystal tracers for full field temperature and velocity measurements, 3d Int. Symp. on Enviromental Hydraulics, CD ROM Proc, pp.00011.1-6, Edts. D. Boyer & R. Rankin, ASU, Tempe 2001.

90. Thermochromic liquid crystals applied for heat transfer research, with J.A. Stasiek, Opto-electronics review, vol. 10(1), pp. 1-10, 2002

91. Observations of flow and turbulence in the nocturnal boundary layer, with P. Monti, H.J. Fernando, M. Princevac, W.C. Chan, E.R. Padyjak, J. Atmospheric Scs. vol. 59(17), 2513-2534, 2002

92. Experimental investigation of vapour bubble growth, with M.-C. Duluc, J. Pakleza, Heat Transfer 2002, vol. 3, pp. 479-485, Elsevier 2002

93. Experimental benchmark for casting problems, with A. Cybulski, T. Michalek, Heat Transfer 2002, vol. 4, pp. 813-819, Elsevier 2002

94. Particle Image Velocimetry and Thermometry in two phase problems in Visualization and Imaging in Transport Phenomena, Annals of New York Academy of Sciences, vol.972, pp.213-222, edts. S. Siedeman & A. Landesberg, NY 2002.

95. Tracking of coherent thermal structures on a heated wall. Part 2: DNS Simulation, with A. Mosyak, G. Hetsroni, Exp. in Fluids 34, pp. 390-396, 2003.

96. Simulation of the water freezing process - numerical benchmarks, with T. Michalek, TASK Quarterly 7: 389-408, 2003.

97. Experimental Methods for Quantitative Analysis of Thermally Driven Flows in Phase Change with Convection, eds. T. Kowalewski, D. Gobin, CISM Course Lecture Notes, vol.449, pp. 167-215, Springer 2004

98. Experimental analysis of vapour bubble growth on heated surface, with K. Pakleza, R. Trzcinski, A. Zachara, Archives of Theromdynamics, 25, pp. 1-12, 2004

99. Visualization of the Flow Structure and Temperature Field in the Region of Mixed Convection, with .E. Fornalik , W. Leiner, J. S. Szmyd, H. Ozoe, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

100. Particle Image Velocimetry (PIV) for Cloud Droplets - Laboratory Investigations, with P. M. Korczyk,S. P. Malinowski,in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

101. Experimental and Numerical Study of Marangoni - Natural Convection, with Le Han Tan, E. Leonardi, T. J. Barber, S. S. Leong, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

102. AHP Setup for Low Laminar Melt Flow Study in Crystal Growth, with M. A. Gonik, J. A. Szymczyk, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

103. Electrospinning of Liquid Jets, with S. Blonski,A. Blasinska, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005

104. Natural Convection for Anomalous Density Variation of Water - Numerical Benchmark, with T. Michalek, B. Sarler, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

105. Thermodynamic Parameters of Vapour Bubble Growth by Image Analysis, with Zachara, J. Pakleza, S. Blonski, R. Trzcinski, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

106. Experiments on Up-slope to Down-slope Transition in an Inclined Box Filled with Water, with K. Dekajlo,H. J. S. Fernando, in Mechanics of 21st Century, ICTAM2004 CD-ROM, edts. W. Gutkowski, T.A. Kowalewski, Springer Verlg. 2005.

107. Optical technique DPIV in measurements of granular material flows, Part 1/3 - plane hoppers, with I. Sielamowicz, S. Blonski, Chem. Eng. Sci., 60, pp. 589-598, 2005

108. Laboratory benchmarks for validating numerical simulation of casting problems, with A. Cybulski, T. Michalek, M. Kowalczyk, Prace IPPT - IFTR Reports, vol. 1/2005, IPPT PAN, Warszawa 2005

109. Natural Convection for Anomalous Density Variation of Water - Numerical Benchmark, with T. Michalek, B. Saler, Progress in Computational Fluid Dynamics 5, pp. 158–170, 2005.

110. Numerical Benchmark Based on Natural Convection of Freezing Water , with T. Michalek, 4th International Conference on Computational Heat and Mass Transfer, vol 1, pp. 7-11, University Cachan 2005

111. Experimental and numerical investigation of natural convection in rectangular inclined enclosure, with K. Dekajlo, w Numerical Heat Transfer 2005,Eurotherm 82, Sept. 13-16, 2005 Krakow, pp. 953-962.

112. Selected laboratory benchmarks for validating numerical simulation of casting problems, with M. Kowalczyk, T. Michalek, A. Cybulski w Numerical Heat Transfer 2005, Eurotherm 82, Sept. 13-16, 2005 Krakow, pp. 589-598.

113. Experiments and modelling of electrospinning process, with S, Blonski, S. Baral, Bulletin Polish Acad. Scs., vol. 53/4, pp. 385-394, 2005.

114. Mixing of cloud and clear air in centimetre scales observed in laboratory, with P. Korczyk, S.P. Malinowski, J. Atmospheric Research 82, pp. 173-182, 2006

115. Digital Particle Image Velocimetry (DPIV) technique in measurements of granular material flows, Part 2 of 3-converging hoppers with I. Sielamowicz, S.Blonski, Chem. Eng. Scs. 61,5307-5317, 2006.

116. Turbulent Flow in a Micro-Channel, with S. Blonski, P. Korczyk, in ASME 4th International Conference on Nanochannels, Microchannels and Minichannels, June 2006, Limerick, CD-ROM Proceedings ISBN 0-7918-3778-5, 96090, p. 1-8, 2006.

117. Investigations of turbulence statistics in the laboratory model of an atmospheric cloud, with P. Korczyk, S. Malinowski in 12th Int. Symposium on Flow Visualization, CD ROM Proceedings, ISBN 0-9533991-8-4, pp. 134.1-10, Goettingen, Optimage Ltd 2006.

118. DPIV technique for granular material flows in silo models, with I. Sielamowicz in 12th Int. Symposium on Flow Visualization, CD ROM Proceedings, ISBN 0-9533991-8-4, pp. 164.1-10, Goettingen, Optimage Ltd 2006.

Dr hab. T. A. Kowalewski Warszawa, January 3, 2007